# Computation of attractor dimension and maximal sums of Lyapunov exponents using polynomial optimization

### v1

@@ 2.4 Sum-of-squares relaxations
In sections 2.2 and 2.3 we have formulated upper bounds on LE sums and on Lyapunov dimension that involve pointwise inequality constraints on functions over sets or, equivalently, pointwise suprema of functions over sets. All of these bounds can be expressed as solutions to optimization problems with pointwise nonnegativity constraints. When all expressions are polynomial, nonnegativity can be strengthened into polynomial SOS conditions. This approach is also the basis of the previous paper [34], and we briefly recall the main concepts here.
In order for a function $S:\mathbb{R}^{n}\to\mathbb{R}$ to satisfy $S(x)\geq 0$ for all $X\in\mathbb{R}^{n}$, it suffices that $S\in\Sigma$, where $\Sigma$ denotes the set of polynomials that are representable as a sum of squares of other polynomials. The set of nonnegative polynomials strictly contains $\Sigma$, and the relationship between these sets has been extensively studied [43]. Often one wants to enforce $S(x)\geq 0$ on a subset of $\mathbb{R}^{n}$ without requiring nonnegativity on all of $\mathbb{R}^{n}$. For instance, one may be interested in a set $\Omega$ with a ‘semiagebraic’ definition in terms of polynomial equalities and inequalities,
@@
where the $g_{i}\in\mathbb{R}[X]$ and $h_{i}\in\mathbb{R}[X]$ are given. Here $\mathbb{R}[X]$ denotes the ring of polynomials in $X$ with real coefficients. There are various ways to define a set $\Sigma^{\Omega}$ of polynomials whose nonnegativity on $\Omega$ is implied by SOS conditions. Here we choose a set of polynomials that is called the quadratic module generated by the polynomials that define $\Omega$,
$\Sigma^{\Omega}=\left\{\sigma_{0}+\sum\limits_{i=1}^{I}\sigma_{i}g_{i}+\sum\limits_{j=1}^{J}\rho_{j}h_{j}~:~\sigma_{i}\in\Sigma\text{ for }i=0,\ldots,I,~\rho_{j}\in\mathbb{R}[X]\text{ for }j=1,\ldots,J\right\},$ (77)
where all $\mathbb{R}[X]$ denotes the set of all polynomial functions of $X$ with real coefficients. Any polynomial in $\Sigma^{\Omega}$ must be nonnegative on $\Omega$ because $\sigma_{0}(X)\geq 0$ at every $X\in\mathbb{R}^{n}$, while each $\sigma_{i}(X)g_{i}(X)\geq 0$ and each $\rho_{j}(X)h_{j}(X)=0$ at every $X\in\Omega$. When $\Omega=\mathbb{R}^{n}$, there are no $g_{i}$ or $\rho_{j}$, so $\Sigma^{\Omega}=\Sigma$. In the preceding paper [34], the condition $S\in\Sigma^{\Omega}$ was expressed in a different but equivalent way: by saying there exist $\sigma_{i}$ and $\rho_{j}$ such that $S-\sum\sigma_{i}g_{i}-\sum\rho_{j}h_{j}$ is SOS.
Computational formulations enforce membership in $\Sigma^{\Omega}$ via a stronger constraint of membership in a finite-dimensional subspace. The simplest choice is to enforce membership in $\Sigma^{\Omega}_{\nu}$, which the truncation of the quadratic module 77 where each $\sigma_{i}$ and $\rho_{j}$ has polynomial degree no larger than $\nu$. This gives a hierarchy of computational formulations, where raising $\nu$ cannot worsen results–and typically improves them–but makes computations more expensive. In practice, it is often useful to choose the spaces for $\sigma_{i}$ and $\rho_{i}$ more carefully. In our computations of section 3, for instance, we impose symmetries and allow different maximum degrees in different variables. For simplicity in the present subsection, however, we write formulations in terms of the simple truncation $\Sigma^{\Omega}_{\nu}$.
The upper bounds on LE sums and Lyapunov dimension formulated in sections 2.2 and 2.3 are stated as, or can be reformulated as, minimizations subject to pointwise inequalities on $V$. When the ODE right-hand side $f(x)$ is polynomial, and when we restrict to auxiliary functions $V$ in some finite-dimensional space of polynomials, then the pointwise inequalities can be enforced using SOS conditions. This gives SOS optimization problems that are computationally tractable when they are not too large, and whose minima give upper bounds on the desired quantities. For the shifted spectrum approach of section 2.3, the SOS relaxations are immediate after choosing a polynomial space for $V$ and enforcing nonnegativity on some set $\Omega$ via membership in $\Sigma^{\Omega}$. For the sphere projection approach of section 2.2, the relevant minmax problems can be expressed as constrained minimizations, after which the SOS relaxations are simple.
@@ 2.4.1 Sphere projection
$\max_{X(0)\in\Omega}\overline{\Phi}\leq\inf_{V\in C^{1}(\Omega)}B\quad\text{s.t.}\quad B-\Phi-F\cdot DV\geq 0~\;\forall X\in\Omega.$ (78)
An SOS relaxation of the right-hand side gives an upper bound, stated in the first part of proposition 2.13 below. Furthermore, this upper bound is an equality if $\Omega$ is forward invariant and its semialgebraic specification 76 has the Archimedean property, as stated in the second part of proposition 2.13 below. The Archimedean property implies compactness of $\Omega$ and is only slightly stronger; any specification of compact $\Omega$ that is not Archimedean can be given this property by adding the polynomial $g_{i}(X)=R^{2}-|X|^{2}$ with a constant $R$ that is large enough to not change $\Omega$ itself. A proof of the second part of proposition 2.13 can be found in [19, Theorem 1]. Briefly, since $\Omega$ is compact and forward invariant, equality in 78 is given by the main theorem of [46], whose applicability on compact manifolds is addressed after (8) in [34]. Then, the Archimedean property guarantees that $\Sigma^{\Omega}$ contains $B-\Phi-F\cdot DV$ for each suboptimal $B$, by Putinar’s Positivstellensatz [41, Lemma 4.1].
@@
Let $\frac{\mathrm{d}}{\mathrm{d}t}X(t)=F(X(t))$ with $F\in\mathbb{R}^{n}[X]$, and let $\Phi\in\mathbb{R}[X]$. Let the dynamics be forward invariant on a semialgebraic set $\Omega\subset\mathbb{R}^{n}$, whose specification defines the quadratic module $\Sigma^{\Omega}$.
@@
    $\max_{X_{0}\in\Omega}\overline{\Phi}\leq\inf_{V\in\mathbb{R}[X]}B\quad\text{s.t.}\quad B-\Phi-F\cdot DV\in\Sigma^{\Omega}.$ (79)
@@
    If $\Omega$ is compact and its semialgebraic specification is Archimedean, then,
@@
All of the upper bounds from section 2.2 can be relaxed into SOS optimization problems by the same reasoning that leads from 5 to 79. Proposition 2.15 below is an SOS relaxation of proposition 2.6 for bounding LE sums. Computations giving such bounds also imply bounds on Lyapunov dimension via proposition 2.7, but potentially sharper bounds on dimension can be computed directly using proposition 2.16, which is an SOS relxation of proposition 2.8. In the proposition statements, $\mathbb{R}^{n}[x]$ denotes the space of polynomials in $x$ taking vector values in $\mathbb{R}^{n}$, and the $x(t)$ domain is a semialgebraic set,
@@
In the following propositions 2.15, 2.16, 2.17 and 2.18, for fixed $k\in\{1,\ldots,n\}$ and polynomial degrees $\nu$ and $\nu^{\prime}$, the computation of the resulting upper bound $C_{i}(k,\nu,\nu^{\prime})$ is an SOS program that can be solved using standard software (see section 3). For certain combinations $(k,\nu,\nu^{\prime})$, there will not exist any $V$ in the chosen space of polynomials that satisfy the SOS constraints, in which case we understand the infimum over bounds to be $\infty$, meaning that we obtain no upper bound. When bounding Lyapunov dimension, choosing $k>d_{L}-1$ is necessary for the constraint set to not be empty. In all cases, raising the polynomial degrees $\nu$ and $\nu^{\prime}$ cannot shrink the constraint set and will often enlarge it.
@@
Let $\frac{\mathrm{d}}{\mathrm{d}t}x(t)=f(x(t))$ with $f\in\mathbb{R}^{n}[x]$, for which a semialgebraic set $\mathcal{B}$ is forward invariant. Let $\Sigma^{\Omega}$ be the quadratic module 77 for the set $\Omega=\mathcal{B}\times\mathbb{S}^{n_{k}}$ whose semialgebraic specification includes the polynomial constraints on $x$ in 81 and the equality $\big|z^{k}\big|-1=0$. Let $\Phi_{k}(x,z^{k})$ and $\ell_{k}(x,z^{k})$ be defined as in 28 and 30, respectively.
@@
    Suppose that each trajectory $x(t)$ in $\mathcal{B}$ is bounded for $t\geq 0$. For each $k\in\{1,\ldots,n\}$, and each pair of polynomial degrees $(\nu,\nu^{\prime})$, the maximum leading LE sum $M_{k}$ among trajectories in $\mathcal{B}$ is bounded above by
@@
    Suppose further that the invariant set $\mathcal{B}$ is compact, and its semialgebraic specification is Archimedean. Then, for each $k$,
@@
The first part is an SOS relaxation of the upper bound 31 on $M_{k}$ for each $k$. Equivalently, it is a particular case of 79 with $X$, $F$ and $\Phi$ as described above in proposition 2.6, where $V$ is further constrained by having degree no larger than $\nu$, and where $\sigma_{i}$ and $\rho_{j}$ in the quadratic module representation 77 have degrees no larger than $\nu^{\prime}$. The second part is the analogous special case of 80. This equality is applicable because the $\mathcal{B}$ is specified as an Archimedean subset of $\mathbb{R}^{n}$, and so adding $|z^{k}|-1=0$ gives an Archimedean specification of $\Omega$ in $\mathbb{R}^{n+n_{k}}$. All polynomial degrees are admitted in 80, which is equivalent to the supremum over $\nu$ and $\nu^{\prime}$ in 84. ∎
@@
Let $\frac{\mathrm{d}}{\mathrm{d}t}x(t)=f(x(t))$ with $f\in\mathbb{R}^{n}[x]$, for which a semialgebraic set $\mathcal{B}$ is forward invariant, and let the integer $k\leq n$ be such that $\sup_{x_{0}\in\mathcal{B}}M_{k+1}(x_{0})<0$. Let $\Sigma^{\Omega}$ be the quadratic module 77 for the set $\Omega=\mathcal{B}\times\mathbb{S}^{n_{k}-1}\times\mathbb{S}^{n_{k+1}-1}$ whose semialgebraic specification includes the polynomial constraints on $x$ in 81, and the equalities $\big|z^{k}\big|-1=0$ and $\big|z^{k+1}\big|-1=0$. Let $\Phi_{k}(x,z^{k})$ and $\ell_{k}(x,z^{k})$ be defined as in 28 and 30, respectively, and likewise for $\Phi_{k+1}(x,z^{k+1})$ and $\ell_{k+1}(x,z^{k+1})$.
@@
    Suppose that each trajectory $x(t)$ in $\mathcal{B}$ is bounded for $t\geq 0$. For each pair of polynomial degrees $(\nu,\nu^{\prime})$, the the maximum Lyapunov dimension $d_{L}$ among trajectories in $\mathcal{B}$ is bounded above by
@@
    $C_{2}(k,\nu,\nu^{\prime},\varepsilon)=\hskip-6.0pt\inf\limits_{\begin{subarray}{c}B\in[0,1]\\ V\in\mathbb{R}[x,z^{k},z^{k+1}]_{\nu}\end{subarray}}\hskip-10.0ptB\quad\text{s.t.}\quad\varepsilon-[\Psi^{B}+f\cdot D_{x}V+\ell_{k}\cdot D_{z^{k}}V+\ell_{k+1}\cdot D_{z^{k+1}}V]\in\Sigma^{\Omega}_{\nu^{\prime}}.$ (86)
@@
    Suppose further that the invariant set $\mathcal{B}$ is compact, and its semialgebraic specification is Archimedean. Let $j$ be the minimum admissible $k$, as in 2.3 for $d_{L}$. Then,
    $d_{L}=j~+\inf_{\begin{subarray}{c}\nu,\nu^{\prime}\in\mathbb{Z}{{}_{\geq}}\\ \varepsilon>0\end{subarray}}C_{2}(j,\nu,\nu^{\prime},\varepsilon).$ (87)
@@
The first part follows from the upper bound 40 on $d_{L}$ because the SOS constraint in 86 with $\varepsilon=0$ implies the nonpositivity constraint on the right-hand side of 40. For the second part, we first establish upper and lower bounds on $C_{2}$ in terms of $d_{L}$. For the lower bound, note that the constraint in 86 implies the constraint in the definition 42 of $B_{\varepsilon}$. This implies $B_{\varepsilon}\leq C_{2}(j,\nu,\nu^{\prime},\varepsilon)$, with which the first inequality in 41 gives
@@
$d_{L}\leq j~+\inf_{\begin{subarray}{c}\nu,\nu^{\prime}\in\mathbb{Z}{{}_{\geq}}\\ \varepsilon>0\end{subarray}}C_{2}(j,\nu,\nu^{\prime},\varepsilon)$ (89)
The infimum will be approached or attained as $\nu,\nu^{\prime}\to\infty$ and $\varepsilon\to 0^{+}$.
For the upper bound on $C_{2}$, note that since $\mathcal{B}$ has an Archimedean specification, so does $\Omega$. Thus we can apply the general SOS upper bounds with equality 80, in particular with $F(X)$ being the right-hand side of the ODE system 37, with $\Phi(X)$ being $\Psi^{B}$, and with $k=j$. This gives
$\sup_{\begin{subarray}{c}x_{0}\in\mathcal{B}~~~~~\\ z^{j}_{0}\in\mathbb{S}^{n_{j}-1}\\ z^{j+1}_{0}\in\mathbb{S}^{n_{j+1}-1}\end{subarray}}\hskip-14.0pt\overline{\Psi}^{B}=\inf_{V\in\mathbb{R}[x,z^{j},z^{j+1}]}\hskip-12.0pt\varepsilon\quad\text{s.t.}\quad\varepsilon-\left[\Psi^{B}+f\cdot D_{x}V+\ell_{k}\cdot D_{z^{j}}V+\ell_{j+1}\cdot D_{z^{j+1}}V\right]\in\Sigma^{\Omega}.$ (90)
Note that the variable called $B$ in 80 is $\varepsilon$ in 90, where $B$ has a different meaning. Let $B$ be such that the constraint holds in expression 47 for $d_{L}$. This constraint is equivalent to both sides of 90 being nonpositive, and nonpositivity of the right-hand infimum in 90 implies that there exists polynomial $V$ satisfying the SOS constraint for any $\varepsilon>0$. Since such $V$ exists for any $B>d_{L}-j$ ,
$d_{L}-j~\geq\inf_{\begin{subarray}{c}B\in[0,1]\\ V\in\mathbb{R}[x,z^{j},z^{j+1}]\end{subarray}}\hskip-8.0ptB\quad\text{s.t.}\quad\varepsilon-\left[\Psi^{B}+f\cdot D_{x}V+\ell_{k}\cdot D_{z^{j}}V+\ell_{j+1}\cdot D_{z^{j+1}}V\right]\in\Sigma^{\Omega}$ (91)
for all $\varepsilon>0$. Restricting the right-hand minimization to $V$ of degree $\nu$ and weights $\sigma_{i}$ and $\rho_{j}$ of degree $\nu^{\prime}$ gives the definition of $C_{2}(j,\nu,\nu^{\prime},\varepsilon)$. Thus, for all $\varepsilon>0$,
@@
(This inequality need not hold with finite $\nu$ and $\nu^{\prime}$.) This upper bound of $d_{L}$, combined with the lower bound of $d_{L}$ 89, established the claim 87 of the second part and completes the proof. ∎
The minimization problems defining $C_{1}$ and $C_{2}$ for each $(k,\nu,\nu^{\prime})$ in 82 of proposition 2.15 are SOS programs. That is, polynomials of finite degree are constrained to be SOS, and these polynomials as well as the optimization objective are affine in the optimization parameters. Here the optimization parameters include $B$ as well as the coefficients of $V$ in whatever basis it is represented. Such SOS programs are computationally tractable so long as the dimension of $X$ and the degree of polynomials in the SOS constraints are not too large. Increasing $\nu$ and $\nu^{\prime}$ gives a hierarchy of SOS programs with increasing computational cost whose minima cannot increase. Typically these minima decrease as $\nu$ and $\nu^{\prime}$ are raised, and the second part of each theorem guarantees convergence to the sharp bound under mild conditions. More generally, a hierarchy of SOS programs can be defined by choosing different sequences of finite-dimensional polynomial spaces, rather than simply truncating by maximum degrees of $\nu$ and $\nu^{\prime}$. This often includes imposing symmetries on each polynomial subspace due to symmetries of $f(x)$, as explained in section 2.5 below.

### v2

@@ 2.4 Sum-of-squares relaxations
The upper bounds on LE sums and Lyapunov dimension formulated in sections 2.2 and 2.3 are stated as, or can be reformulated as, minimizations subject to pointwise inequalities on $V$. When the ODE right-hand side $f(x)$ is polynomial, and when we restrict to auxiliary functions $V$ in some finite-dimensional space of polynomials, then the pointwise inequalities can be enforced using SOS conditions. This gives SOS optimization problems that are computationally tractable when the polynomial degrees are not too large, and whose minima give upper bounds on the desired quantities. For the shifted spectrum approach of section 2.3, the SOS relaxations are immediate after choosing a polynomial space for $V$ and enforcing nonnegativity on some set $\Omega$ via membership in $\Sigma^{\Omega}$. For the sphere projection approach of section 2.2, the relevant minmax problems can be expressed as constrained minimizations, after which the SOS relaxations are simple. We review the necessary basics of SOS polynomials here.
In order for a polynomial $S:\mathbb{R}^{n}\to\mathbb{R}$ to satisfy $S(x)\geq 0$ for all $X\in\mathbb{R}^{n}$, it suffices that $S\in\Sigma$, where $\Sigma$ denotes the set of polynomials that are representable as a sum of squares of other polynomials. The set of nonnegative polynomials strictly contains $\Sigma$, and the relationship between these sets has been extensively studied [46]. Often one wants to enforce $S(x)\geq 0$ on a subset of $\mathbb{R}^{n}$ without requiring nonnegativity on all of $\mathbb{R}^{n}$. For instance, one may be interested in a set $\Omega$ with a ‘semiagebraic’ definition in terms of polynomial equalities and inequalities,
@@
where the $g_{i}\in\mathbb{R}[X]$ and $h_{i}\in\mathbb{R}[X]$ are given. Here $\mathbb{R}[X]$ denotes the ring of polynomials in $X$ with real coefficients, and $\mathbb{R}^{n}[X]$ denotes the vectors of such polynomials taking values in $\mathbb{R}^{n}$. There are various ways to define a set $\Sigma^{\Omega}$ of polynomials whose nonnegativity on $\Omega$ is implied by SOS conditions. Here we choose $\Sigma^{\Omega}$ to be the quadratic module generated by the polynomials that define $\Omega$,
$\Sigma^{\Omega}=\left\{\sigma_{0}+\sum\limits_{i=1}^{I}\sigma_{i}g_{i}+\sum\limits_{j=1}^{J}\rho_{j}h_{j}~:~\sigma_{i}\in\Sigma\text{ for }i=0,\ldots,I,~\rho_{j}\in\mathbb{R}[X]\text{ for }j=1,\ldots,J\right\}.$ (74)
Any polynomial in $\Sigma^{\Omega}$ must be nonnegative on $\Omega$ because $\sigma_{0}(X)\geq 0$ at every $X\in\mathbb{R}^{n}$, while each $\sigma_{i}(X)g_{i}(X)\geq 0$ and each $\rho_{j}(X)h_{j}(X)=0$ at every $X\in\Omega$. When $\Omega=\mathbb{R}^{n}$, there are no $g_{i}$ or $\rho_{j}$, so $\Sigma^{\Omega}=\Sigma$. In the preceding paper [34], the condition $S\in\Sigma^{\Omega}$ was expressed in a different but equivalent way: by saying there exist $\sigma_{i}$ and $\rho_{j}$ such that $S-\sum\sigma_{i}g_{i}-\sum\rho_{j}h_{j}$ is SOS.
Computational formulations enforce membership not only in $\Sigma^{\Omega}$ but in a finite-dimensional subset. The simplest choice is to enforce membership in $\Sigma^{\Omega}_{\nu}$, which is the truncation of the quadratic module 74 where each $\sigma_{i}$ and $\rho_{j}$ has polynomial degree no larger than $\nu$, for some $\nu$ in the set $\mathbb{Z}_{\geq}$ of nonnegative integers. This gives a hierarchy of computational formulations, in which raising $\nu$ cannot worsen results—and typically improves them—but makes computations more expensive. In practice, it is often useful to choose the spaces for $\sigma_{i}$ and $\rho_{i}$ more carefully. In our computations of section 3, for instance, we impose symmetries and allow different maximum degrees in different variables. For simplicity in the present subsection, however, we write formulations in terms of the degree-$\nu$ truncation $\Sigma^{\Omega}_{\nu}$.
@@ 2.4.1 Sphere projection
$\max_{X_{0}\in\Omega}\overline{\Phi}\leq\inf_{V\in C^{1}(\Omega)}B\quad\text{s.t.}\quad B-\Phi-F\cdot DV\geq 0~\;\forall X\in\Omega.$ (75)
An SOS relaxation of the right-hand side gives an upper bound, stated in the first part of proposition 2.14 below. Furthermore, this upper bound is an equality if $\Omega$ is forward invariant and its semialgebraic specification 73 has the Archimedean property, as stated in the second part of proposition 2.14 below. The Archimedean property implies compactness of $\Omega$ and is only slightly stronger; any semialgebraic specification of compact $\Omega$ that is not Archimedean can be given this property by adding the polynomial $g_{i}(X)=R^{2}-|X|^{2}$ with a constant $R$ that is large enough to not change $\Omega$ itself. A proof of the second part of proposition 2.14 can be found in [20, Theorem 1]. Briefly, since $\Omega$ is compact and forward invariant, equality in 75 is given by the main theorem of [50], whose applicability on compact manifolds is addressed after (8) in [34]. Then, for each suboptimal $B$, there exists a polynomial $V$ such that $B-\Phi-F\cdot DV>0$ on $\Omega$, and Putinar’s Positivstellensatz [43, Lemma 4.1] guarantees that $B-\Phi-F\cdot DV$ belongs to $\Sigma^{\Omega}$.
@@
Let $X(t)$ evolve by the ODE 1, where $F\in\mathbb{R}^{n}[X]$. Let $\Omega\subset\mathbb{R}^{n}$ be a semialgebraic set that is forward invariant, and whose quadratic module is denoted $\Sigma^{\Omega}$. Let $\Phi\in\mathbb{R}[X]$.
@@
    $\sup_{X_{0}\in\Omega}\overline{\Phi}\leq\inf_{V\in\mathbb{R}[X]}B\quad\text{s.t.}\quad B-\Phi-F\cdot DV\in\Sigma^{\Omega}.$ (76)
@@
    If $\Omega$ is compact and its semialgebraic specification is Archimedean, then
@@
All of the upper bounds from section 2.2 can be relaxed into SOS optimization problems by the same reasoning that leads from 5 to 75 to 76. Proposition 2.16 below is an SOS relaxation of proposition 2.6 for bounding LE sums. Computations giving such bounds also imply bounds on Lyapunov dimension via proposition 2.7, but potentially sharper bounds on dimension can be computed directly using proposition 2.17, which is an SOS relaxation of proposition 2.8. In the proposition statements, the domain $\mathcal{B}$ is assumed to have the semialgebraic specification
@@
In the following propositions 2.16, 2.17, 2.18 and 2.19, for fixed $k\in\{1,\ldots,n\}$ and polynomial degrees $\nu$ and $\nu^{\prime}$, the resulting upper bound $C_{i}(k,\nu,\nu^{\prime})$ can be computed by solving SOS programs using standard software (see section 3). For certain combinations $(k,\nu,\nu^{\prime})$, there will not exist any $V$ in the chosen space of polynomials that satisfy the SOS constraints, in which case we understand the infimum over bounds to be $\infty$, meaning that we obtain no upper bound. In propositions 2.16 and 2.17, the values $C_{1}$ and $C_{2}$ are found by solving a single SOS program in which the bound $B$ is minimized. In propositions 2.18 and 2.19, the value $B$ must be fixed in each SOS program, which determines feasibility or infeasibility of the constraints with this $B$ value. In these cases the minimum feasible $B$ is found by repeatedly solving SOS programs to find $C_{3}$ and $C_{4}$.
@@
Let $x(t)$ evolve by the ODE 6, where $f\in\mathbb{R}^{n}[X]$ and $\mathcal{B}\subset\mathbb{R}^{n}$ is a semialgebraic set that is forward invariant. Let $\Sigma^{\Omega}$ be the quadratic module 74 for the set $\Omega=\mathcal{B}\times\mathbb{S}^{C_{n}^{k}-1}$ whose semialgebraic specification includes the polynomial constraints on $x$ in 78 and the equality $\big|z^{k}\big|-1=0$. Let $\Phi_{k}(x,z^{k})$ and $\ell_{k}(x,z^{k})$ be as defined in 28 and 30, respectively.
@@
    Suppose each trajectory $x(t)$ in $\mathcal{B}$ is bounded for $t\geq 0$. For each $k\in\{1,\ldots,n\}$, and each pair of polynomial degrees $(\nu,\nu^{\prime})$, the maximum leading LE sum $M_{k}$ among trajectories in $\mathcal{B}$ is bounded above by
@@
    Suppose $\mathcal{B}$ is compact and its semialgebraic specification is Archimedean. Then, for each $k$,
@@
The first part is an SOS relaxation of the upper bound 31 on $M_{k}$ for each $k$. Equivalently, it is a particular case of 76 with $X$, $F$ and $\Phi$ as described above in proposition 2.6, where $V$ is further constrained by having degree no larger than $\nu$, and where $\sigma_{i}$ and $\rho_{j}$ in the quadratic module representation 74 have degrees no larger than $\nu^{\prime}$. The second part is the analogous special case of 77. This equality is applicable because $\mathcal{B}$ is specified as an Archimedean subset of $\mathbb{R}^{n}$, and so adding $|z^{k}|-1=0$ gives an Archimedean specification of $\Omega$ in $\mathbb{R}^{n}\times\mathbb{R}^{C_{n}^{k}}$. All polynomial degrees are admitted in 77, which is equivalent to the supremum over $\nu$ and $\nu^{\prime}$ in 81. ∎
@@
Let $x(t)$ evolve by the ODE 6, where $f\in\mathbb{R}^{n}[X]$ and $\mathcal{B}\subset\mathbb{R}^{n}$ is a semialgebraic set that is forward invariant. Let the positive integer $k<n$ be such that $\sup_{x_{0}\in\mathcal{B}}M_{k+1}(x_{0})<0$. Let $\Sigma^{\Omega}$ be the quadratic module 74 for the set $\Omega=\mathcal{B}\times\mathbb{S}^{C_{n}^{k}-1}\times\mathbb{S}^{C_{n}^{k+1}-1}$, whose semialgebraic specification includes the polynomial constraints on $x$ in 78 and the equalities $\big|z^{k}\big|-1=0$ and $\big|z^{k+1}\big|-1=0$. Let $\Phi_{k}(x,z^{k})$ and $\ell_{k}(x,z^{k})$ be defined as in 28 and 30, respectively, and likewise for $\Phi_{k+1}(x,z^{k+1})$ and $\ell_{k+1}(x,z^{k+1})$.
@@
    Suppose each trajectory $x(t)$ in $\mathcal{B}$ is bounded for $t\geq 0$. For each pair of polynomial degrees $(\nu,\nu^{\prime})$, the maximum Lyapunov dimension $d_{L}$ among trajectories in $\mathcal{B}$ is bounded above by
@@
    $C_{2}(k,\nu,\nu^{\prime},\varepsilon)=\hskip-6.0pt\inf\limits_{\begin{subarray}{c}B\in[0,1]\\ V\in\mathbb{R}[x,z^{k},z^{k+1}]_{\nu}\end{subarray}}\hskip-10.0ptB\quad\text{s.t.}\quad\varepsilon-[\Psi^{B}+f\cdot D_{x}V+\ell_{k}\cdot D_{z^{k}}V+\ell_{k+1}\cdot D_{z^{k+1}}V]\in\Sigma^{\Omega}_{\nu^{\prime}},$ (83)
    with $C_{2}=+\infty$ if the constraint set is empty.
@@
    Suppose $\mathcal{B}$ is compact and its semialgebraic specification is Archimedean. Let $j$ be the minimum admissible $k$, as in 2.3 for $d_{L}$. Then,
    $d_{L}=j~+\inf_{\begin{subarray}{c}\nu,\nu^{\prime}\in\mathbb{Z}{{}_{\geq}}\\ \varepsilon\geq 0\end{subarray}}C_{2}(j,\nu,\nu^{\prime},\varepsilon).$ (84)
@@
The first part follows from the upper bound 40 on $d_{L}$ because the SOS constraint in 83 with $\varepsilon=0$ implies the nonpositivity constraint on the right-hand side of 40. For the second part, we first establish upper and lower bounds on $C_{2}$ in terms of $d_{L}$. For the lower bound, note that the constraint in 83 implies the constraint in the definition 42 of $B_{\varepsilon}$. This implies $B_{\varepsilon}\leq C_{2}(j,\nu,\nu^{\prime},\varepsilon)$, with which the first inequality in 41 gives
@@
$d_{L}\leq j~+\inf_{\begin{subarray}{c}\nu,\nu^{\prime}\in\mathbb{Z}{{}_{\geq}}\\ \varepsilon>0\end{subarray}}C_{2}(j,\nu,\nu^{\prime},\varepsilon),$ (86)
and this inequality remains true when the constraints allow $\varepsilon=0$, due to 82.
For the upper bound on $C_{2}$, note that since $\mathcal{B}$ has an Archimedean specification, so does $\Omega$. Thus we can apply the general SOS equality 77, in particular with $F(X)$ being the right-hand side of the ODE system 37, with $\Phi(X)$ being $\Psi^{B}$, and fixing $k=j$. This gives
$\sup_{\begin{subarray}{c}x_{0}\in\mathcal{B}~~~~~\\ z^{j}_{0}\in\mathbb{S}^{C_{n}^{j}-1}\\ z^{j+1}_{0}\in\mathbb{S}^{C_{n}^{j+1}-1}\end{subarray}}\hskip-14.0pt\overline{\Psi}^{B}=\inf_{V\in\mathbb{R}[x,z^{j},z^{j+1}]}\hskip-12.0pt\varepsilon\quad\text{s.t.}\quad\varepsilon-\left[\Psi^{B}+f\cdot D_{x}V+\ell_{j}\cdot D_{z^{j}}V+\ell_{j+1}\cdot D_{z^{j+1}}V\right]\in\Sigma^{\Omega}.$ (87)
Note that the variable called $B$ in 77 is $\varepsilon$ in 87, and here $B$ has a different meaning. Let $B$ be such that the constraint holds in expression 47 for $d_{L}$. This constraint is equivalent to both sides of 87 being nonpositive, and nonpositivity of the right-hand infimum in 87 implies that there exists polynomial $V$ satisfying the SOS constraint for any $\varepsilon>0$. Since such $V$ exists for any $B>d_{L}-j$ ,
$d_{L}-j~\geq\inf_{\begin{subarray}{c}B\in[0,1]\\ V\in\mathbb{R}[x,z^{j},z^{j+1}]\end{subarray}}\hskip-8.0ptB\quad\text{s.t.}\quad\varepsilon-\left[\Psi^{B}+f\cdot D_{x}V+\ell_{j}\cdot D_{z^{j}}V+\ell_{j+1}\cdot D_{z^{j+1}}V\right]\in\Sigma^{\Omega}$ (88)
for all $\varepsilon>0$. Restricting the right-hand minimization to $V$ of degree $\nu$ and to $\sigma_{i}$ and $\rho_{j}$ of degree $\nu^{\prime}$ gives the definition of $C_{2}(j,\nu,\nu^{\prime},\varepsilon)$. Thus, for all $\varepsilon>0$,
@@
(This inequality need not hold with any finite $\nu$ and $\nu^{\prime}$.) The upper bound 89 on $C_{2}$, combined with the lower bound 86, establishes the claim 84 of the second part and completes the proof. ∎
For each fixed $(k,\nu,\nu^{\prime})$, the minimization problems defining $C_{1}$ in 80 and $C_{2}$ in 83 are SOS programs. That is, polynomials of finite degree are constrained to be SOS, and these polynomials, as well as the optimization objective, are affine in the optimization parameters. Here the optimization parameters are $B$ and the coefficients of $V$ (which depend on the basis in which $V$ is represented). Such SOS programs are computationally tractable so long as the number of variables and the polynomial degrees in the SOS constraints are not too large. Increasing $\nu$ and $\nu^{\prime}$ gives a hierarchy of SOS programs with increasing computational cost whose minima cannot increase. Typically these minima decrease as $\nu$ and $\nu^{\prime}$ are raised, and the second part of each proposition guarantees convergence to the sharp bound under mild conditions.